Calculate the power in a 1000 Ω electric heater if the current through it is (a) 0.5 A & (b) 1A.

providing the given, required, formula,

solution/substitution of data


1
Expert's answer
2022-03-30T13:51:17-0400

Answer

a) power

"P=\frac{I^2}{R}\\=\frac{0.5*0.5}{1000}\\=0.00025A"

b) "P=\frac{I^2}{R}\\=\frac{1*1}{1000}\\=0.001A"